# LLMs as Planning Modelers: A Survey for Leveraging Large Language Models to Construct Automated Planning Models

### v1

@@ 27 FEEDBACK:
29 All necessary objects are included based on the problem description. So: No.
@@
31 All objects included are relevant to the problem. Hence: No.
@@
35 All predicates used in the initial state are relevant and correctly applied. Thus: No.
@@
37 The initial state is missing the predicate for the red block being clear. Since the red block is covered by
38 the blue block, it should not be clear. Therefore: Yes.
@@
40 The goal state accurately reflects the desired outcome of having the red block on top of the green block. So: No.
@@
44 The predicates used do not require symmetry as they are directional in nature. So: No.
45
46 My concrete suggestions are the following:
47 - Add the predicate to indicate that the red block is not clear:
48 - (clear red_block) should be removed from the initial state since the red block is covered by the blue block.
49
50 Final output should reflect this change in the initial state:
@@
52 (clear blue_block): blue block is clear
53 (clear yellow_block): yellow block is clear
54 (clear green_block): green block is clear
55 ‘‘‘
56
57 Overall, the feedback is: Yes, the initial state needs to reflect that the red block is not clear.
@@
Figure 7: L2P formatted predicate output from LLM (action-by-action algorithm) of Logistics domain.
Appendix C Paper Overview
This section contains an overview of the core framework papers found within this survey. The table provides a structured distinction between the papers based on their methodological approaches and specific areas of focus, such as what aspect are they modeling, prompting style/weight, and feedback mechanism. Additionally, it highlights papers already covered by the L2P library, ensuring users have quick access to resources that have been implemented or analyzed in detail. These covered papers will be included in our GitHub repository for easy reference. The distinction in the table also aims to emphasize gaps in the literature, such as under explored techniques or novel applications of NL-PDDL through LLMs, paving the way for future research. To maintain relevance, we plan to keep an updated installment of all NL-PDDL works leveraging LLMs, ensuring this survey evolves alongside the field.
Framework Task Domain NHI Prompt Feedback
(Collins et al. 2022) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) None
(Xie et al. 2023) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) None
(Lyu et al. 2023) $\checkmark$ $\times$ $\checkmark$ Light (CoT) None
(Grover and Mohan 2024) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) Env.
(Lin et al. 2023) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) LLM+Env.
(Liu et al. 2023a) $\checkmark$ $\times$ $\checkmark$ Heavy (Few-shot) None
(Birr et al. 2024) $\checkmark$ $\times$ $\checkmark$ Light (CoT) External tool
(Agarwal and Sreepathy 2024) $\checkmark$ $\times$ $\times$ Medium None
(Dagan, Keller, and Lascarides 2023) $\checkmark$ $\times$ $\checkmark$ Medium LLM + Env.
(Zhang et al. 2024a) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) LLM
(Liu et al. 2024c) $\checkmark$ $\times$ $\times$ Extended Scene Graph Env.
(Singh et al. 2024) $\checkmark$ $\times$ $\times$ Light LLM+Deviation
(Izquierdo-Badiola et al. 2024) $\checkmark$ $\times$ $\checkmark$ Light LLM
(Singh, Traum, and Thomason 2024) $\checkmark$ $\times$ $\checkmark$ Medium (One-shot+) LLM Helper Agent
(Zhang et al. 2024d) $\checkmark$ $\times$ $\checkmark$ Light FastDownward+External Tool
(Oates et al. 2024) $\times$ $\checkmark$ $\times$ Medium (Few-shot, in-context) Human
(Zhang et al. 2024b) $\times$ $\checkmark$ $\times$ Medium (Few-shot) None
(Guan et al. 2023) $\times$ $\checkmark$ $\times$ Light Human
(Huang, Lipovetzky, and Cohn 2024) $\times$ $\checkmark$ $\checkmark$ Light Sentence Encoder Filter
(Wong et al. 2023) $\times$ $\checkmark$ $\checkmark$ Medium (Few-shot) None
(Liu et al. 2024a) $\times$ $\checkmark$ $\times$ Heavy Human
(Ding et al. 2023a) $\times$ $\checkmark$ $\checkmark$ Light LLM
(Chen et al. 2024) $\times$ $\checkmark$ $\checkmark$ Medium Env.
(Kelly et al. 2023) $\checkmark$ $\checkmark$ $\checkmark$ Medium (One-shot+) External tool
(Smirnov et al. 2024) $\checkmark$ $\checkmark$ $\checkmark$ Light LLM
(Liu et al. 2024b) $\checkmark$ $\checkmark$ $\checkmark$ Medium (One-shot+) None
(Zhou et al. 2023) $\checkmark$ $\checkmark$ $\checkmark$ Heavy (Few-shot) None
(Ye et al. 2024) $\checkmark$ $\checkmark$ $\times$ Heavy Human
(Han et al. 2024) $\checkmark$ $\checkmark$ $\times$ Light Human
(Gestrin, Kuhlmann, and Seipp 2024) $\checkmark$ $\checkmark$ $\checkmark$ Heavy (Few-shot) LLM + Human
(da Silva et al. 2024) $\checkmark$ $\checkmark$ $\checkmark$ Light LLM
(Mahdavi et al. 2024) $\checkmark$ $\checkmark$ $\checkmark$ Light LLM + Env.
(Sakib and Sun 2024) $\checkmark$ $\checkmark$ $\checkmark$ Light None
(Ying et al. 2023) $\checkmark$ $\checkmark$ $\checkmark$ Medium (Few-shot) Syntax rejection filter
Table 1: Summary of frameworks in this survey. NHI = No Human Intervention. Papers reconstructed by L2P

### v2

@@ 29 My feedback on the provided PDDL problem file is as follows:
32 All necessary blocks are included. Therefore: No.
33
@@
35 All objects are relevant to the problem. Hence: No.
36
@@
41 All predicates used are relevant and correctly applied. Thus: No.
42
@@
44 The initial state is missing the predicate for the red block being not clear, as it is covered by the blue block. This should be added:
45 - (clear red_block) should be false, but it is not explicitly stated. Hence: Yes.
46
@@
48 The goal state is correctly defined as having the red block on top of the green block. So: No.
49
@@
54 The predicates used do not require symmetry in this context. Hence: No.
55
56 In summary, the main issue is the missing predicate regarding the clarity of the red block. It should be explicitly stated that the red block is not clear due to the blue block being on top of it.
57
58 To improve the initial state, you should add:
59 - (clear red_block) should be false, or alternatively, you can add:
60 - (not (clear red_block)) to indicate that the red block is not clear.
@@ 58 To improve the initial state, you should add:
## ACTIONS
{’name’: ’load_truck’, ’parameters’: OrderedDict([(’?p’, ’package’), (’?t’, ’truck’), (’?l’, ’location’)]),
[…]
{’name’: ’unload_truck’, ’parameters’: OrderedDict([(’?p’, ’package’), (’?t’, ’truck’), (’?l’, ’location’)]),
[…]
{’name’: ’load_airplane’, ’parameters’: OrderedDict([(’?p’, ’package’), (’?a’, ’plane’)]),
[…]
{’name’: ’unload_airplane’, ’parameters’: OrderedDict([(’?p’, ’package’), (’?a’, ’plane’), (’?l’, ’location’)]),
[…]
{’name’: ’drive_truck’,
    ’parameters’: OrderedDict([(’?t’, ’truck’), (’?l1’, ’location’), (’?l2’, ’location’), (’?c’, ’city’)]),
[…]
{’name’: ’fly_airplane’,
    ’parameters’: OrderedDict([(’?a’, ’plane’), (’?l1’, ’location’), (’?l2’, ’location’), (’?c’, ’city’)]),
[…]
Figure 8: L2P formatted predicate and actions outputted from LLM (gpt-4o-mini) via ‘action-by-action‘ algorithm (Guan et al., 2023) on Logistics domain.